I really need help on this one.
4/3x^2 - 2x + 3/4 = 0
I have to use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.

Answer by Prithwis(166)

Discriminant = (b^2 - 4ac) for ax^2+bx+c
Here a = 4/3, b=-2 and c = 3/4
b^2 - 4ac = 4 - 4 = 0
If b^2 - 4ac =0, there is only one real solution.
If b^2 - 4ac > 0, both the roots are real.
If b^2 - 4ac < 0, both the roots are non-real complex.
Here the discriminant is 0; so there is only one real solution of the given equation.